OpenStudy (anonymous):

A bucket of mass 2.50kg is whirled in a vertical circle of radius 1.45m . At the lowest point of its motion the tension in the rope supporting the bucket is 33.0 N. (a) Find the speed of the bucket (b) How fast must the bucket move at the top of the circle so that the rope does not go slack? we did not finish going over this circular motion and gravitational chapter, but I would like to get ahead on work. What formulas will I need for parts a and b? :O

OpenStudy (anonymous):

Don't just look for the formula. try to grab the concept. Here you will need a centrifugal force that acts against the tension which is \[mv ^{^{2}}/r\] and tension and centrifugal force will be equal or else the string will become slack or break

OpenStudy (anonymous):

Ok, so for part b) I did: Energy Fr=(mv^2/r) mg=(mv^2/r) v^2=gr v=square root (gr) \[\sqrt{9.8m/s^2*1.45m}\] =3.77 m/s and that answer was right. But for part a I keep getting the wrong answer. Do I set the formula like I did in part b? Wouldn't the mass cancel out if I did mg=(mv^2/r)?

OpenStudy (anonymous):

For the first part resultant of mv^2/r and mg should be equal to tension. Sorry I cannot help more than this. I m using an iPad the website gets really messy on iPad.